Algebra Word problem:
A sandwich is served on a 18 inch bun.
Jenny, Demi, and Sarah agree to share the sandwich. Due to different appetites
Jenny cuts the sandwich in a way that Demi gets half of what Jenny gets and Sarah gets three-fourths of what Jenny gets.
Find the length of each person's sandwich.
Thanks.

Answer by rothauserc(4718)

A sandwich is served on a 18 inch bun.
Jenny, Demi, and Sarah agree to share the sandwich. Due to different appetites
Jenny cuts the sandwich in a way that Demi gets half of what Jenny gets and Sarah gets three-fourths of what Jenny gets.
Find the length of each person's sandwich.
let x be Jenny's length of sandwhich, then
x + x(1/2) + (3/4)x = 18
(4/4)x + (2/4)x + (3/4)x = 18 (convert all fractions to a common denominator)
(9/4)x = 18 (combine fractions)
9x = 72 (multiply both sides of = by 4)
x = 8
Jenny's sandwich is 8 inches
Demi's sandwich is 4 inches (1/2 of Jenny's)
Sarah's sandwich is 6 inches (3/4 of Jenny's)